Deming & Sons manufactures four grades of lubricant, W-10, W-20, W-30, and W-40, from a joint process. Additional information follows.

If Processed Further
Product Units Produced Sales Value at Split-Off Additional Costs Sales Values
W-10 53,200 $ 316,000 $ 34,100 $ 346,000
W-20 38,000 272,000 27,200 316,000
W-30 30,400 184,000 18,400 228,000
W-40 30,400 140,000 11,400 152,000
152,000 $ 912,000 $ 91,100 $ 1,042,000

Required:

Assuming that total joint costs of $410,400 were allocated using the physical quantities method, what joint costs were allocated to each product? (Do not round intermediate calculations.)
Answer #1

Joint Cost allocation

Product Units produced Weight Joint cost Allocated joint cost
W-10 53200 35% 410400 143640
W-20 38000 25% 410400 102600
W-30 30400 20% 410400 82080
W-40 30400 20% 410400 82080
Total 152000 100% 410400
Answer #2

410,400*53,200/152,000= 143,640

410,000*38,000/152,000= 102,600

410,400*30,400/152,000= 82,080